OpenStudy (anonymous):

find the general solution to 2x(y+1)-yy'=0

OpenStudy (anonymous):

its a differential equation

OpenStudy (anonymous):

but im stuck in the middle

OpenStudy (anonymous):

wait i got it

OpenStudy (anonymous):

but if u have \[\int\limits (y(1/(y+1)) \] why do u subtract when u seperate

OpenStudy (anonymous):

\[\int\limits y -\int\limits 1/(1+y)\]

OpenStudy (anonymous):

why is it subtraction

OpenStudy (anonymous):

wait im wrong

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2x(y+1)=yy' 2x(y+1)/y = y' 2x (1 + 1/y)=y' 2x = y'/(1 + 1/y) 2x dx= dy/(1+1/y)

OpenStudy (anonymous):

ok just answer me this

OpenStudy (anonymous):

if u want to seperate integrals, do u have to subtract?